OpenStudy (anonymous):

For which of these functions f(x) does the limit f(x) = 2 as x -> -∞? a) (x-2)/(3x-5) b) (2x)/radical(x-2) c) (2x^2-6x+1)/(1+x^2) d) (2x-1)/(x^2+1) e) none of the above

OpenStudy (anonymous):

c does

OpenStudy (anonymous):

since numerator and denominator both have degree 2, take the ratio of the leading coefficients and see that you get 2 as a limit

OpenStudy (across):

I will say C:\[\lim_{x\to\infty}\frac{2x^2-6x+1}{1+x^2}=\lim_{x\to\infty}\frac{4x-6}{2x}=\lim_{x\to\infty}\frac{4}{2}=2.\]

hero (hero):

2x^2/x^2 = 2

OpenStudy (anonymous):

Thanks so much l'hopital's rule makes so much sense O.O

hero (hero):

You didn't need to use l'hopital rule for this

OpenStudy (anonymous):

It makes the most sense the way across used it though, it just seems easier, then remembering the exponent rules